The form x2 + y2 + z2 and x2 - y2 - z2 are equivalent. True or false
1
Expert's answer
2019-06-27T13:42:07-0400

It is true because there exists a nonsingular linear transformation

"X = BY,"


where X is a matrix of the first quadratic form x2+y2+z2 ,

and

"X=\begin{pmatrix}

1 & 0 & 0\\

0& 1 & 0\\

0& 0 & 1

\end{pmatrix}"


Y is a matrix of the second quadratic form x2-y2-z2

and


"Y=\begin{pmatrix}

1 & 0 & 0\\

0& -1 & 0\\

0& 0 &- 1

\end{pmatrix}"

and B is a matrix of the nonsingular linear transformation and


"B=\begin{pmatrix}

1 & 0 & 0\\

0& -1 & 0\\

0& 0 &- 1

\end{pmatrix}"

and

"Det(B)=1\neq0"